OpenStudy (anonymous):

I keep on getting the answer wrong when I multiply can anybody work out this for me 80* 2.36 Every timei multiply I always get the Wong answer

OpenStudy (anonymous):

The answer is 188.8, what do you get?

OpenStudy (anonymous):

in this case it's not that hard to do it from your head \[80 * 2 = 8*2*10\]=160 \[80*0.3 = 8*10*\frac{ 3 }{ 10 } = 8*3*\frac{ 10 }{ 10 } = 8*3\]=24 \[80*0.06=8*0.6=\frac{ 8*6 }{ 10 }=\frac{ 48 }{ 10 }=4.8\] add everything up 160+24+4.8 = 188.8